What should be the price of a 3-month futures on the USD/EUR exchange rate when the USD interest rate is 6% (continuously compounded), the EUR interest rate is 4% and the spot exchange rate is EUR 1 = USD1.20?
Answer #1

The future rateF0=spot*e^(usd libor-eur libor)*t
Here t=3months=3/12
=1.2*e^(6%-4%)*(3/12)
=1.206